# SplatCo: Structure-View Collaborative Gaussian Splatting for Detail-Preserving Rendering of Large-Scale Unbounded Scenes

### v3

@@ Abstract
Neural Radiance Fields (NeRFs) have achieved impressive results in novel view synthesis but are less suited for large-scale scene reconstruction due to their reliance on dense per-ray sampling and deep MLP inference, which limits scalability and efficiency. In contrast, 3D Gaussian Splatting (3DGS) offers a more efficient alternative to computationally intensive volume rendering, enabling faster training and real-time rendering. Although recent efforts have extended 3DGS to large-scale settings, these methods often struggle to balance global structural coherence with local detail fidelity, and suffer from view-dependent artifacts and geometric inconsistencies, especially in sparsely observed regions. To address these challenges, we present SplatCo, a structure–view collaborative Gaussian splatting framework for high-fidelity rendering of complex outdoor environments. SplatCo builds upon two novel components: (1) a cross-structure collaboration module that combines global tri-plane representations, which capture coarse scene layouts, with local context grid features that represent fine surface details. This fusion is achieved through a novel hierarchical compensation strategy, ensuring both global consistency and local detail preservation; and (2) a cross-view assisted training strategy that enhances multi-view consistency by synchronizing gradient updates across viewpoints, applying visibility-aware densification, and pruning overfitted or inaccurate Gaussians based on structural consistency. Through joint optimization of structural representation and multi-view coherence, SplatCo effectively reconstructs fine-grained geometric structures and complex textures in large-scale scenes. Comprehensive evaluations on 13 diverse large-scale scenes, including Mill19, MatrixCity, Tanks & Temples, WHU, and custom aerial captures, demonstrate that SplatCo consistently achieves higher reconstruction quality than state-of-the-art methods, with PSNR improvements of 1–2 dB and SSIM gains of 0.1 to 0.2. These results establish a new benchmark for high-fidelity rendering of large-scale unbounded scenes. Code and additional information are available at https://github.com/SCUT-BIP-Lab/SplatCo.
I Introduction
@@
Neural scene representations have significantly advanced novel view synthesis and 3D reconstruction. Among them, Neural Radiance Fields (NeRF) [1] achieve impressive photorealistic rendering of objects and small-scale scenes from multi‑view captures, supporting applications in augmented/virtual reality [2], autonomous driving [3], digital twins [4], telepresence [5], and cultural heritage preservation [6].
Applying NeRFs to large-scale outdoor and urban environments remains difficult due to both computational and representational challenges. These scenes are typically unbounded, which introduces numerical instability in volume rendering. NeRF’s reliance on dense per-ray sampling and deep MLP inference leads to high computational cost, limiting scalability. In addition, NeRFs have limited capacity to model extended spatial regions, often resulting in blurred geometry and textures. Per-view training can also cause single-view overfitting, where reconstructions conform too closely to individual camera rays. This issue becomes more pronounced in sparsely observed areas and reduces consistency across views. Finally, there is a fundamental trade-off between global and local representations: global factorizations capture coarse structure but miss surface detail, while local grids better preserve detail but may become misaligned without a consistent global frame.
@@
To overcome these limitations, 3D Gaussian Splatting (3DGS) [7] has emerged as an efficient alternative. It represents a scene as a sparse collection of anisotropic 3D Gaussians whose positions, shapes, orientations, and colors are optimized directly from images. These Gaussians are rendered using a tile-based GPU splatting pipeline, enabling high-quality, real‑time rendering.
While 3DGS addresses many of the efficiency and scalability issues inherent in NeRF, extending it to large-scale environments introduces new challenges related to structure preservation, detail fidelity, and multi-view consistency. Several strategies have been proposed to scale 3DGS to campus- or city-scale scenes. Tri‑plane factorization [8, 9] compresses 3D structure into three orthogonal feature planes, which efficiently capture global layout. However, this representation tends to blur or omit fine surface details, particularly along curved or oblique regions not aligned with the coordinate axes. Context‑grid augmentations [10, 11] divide the scene into local voxel grids to recover high‑frequency detail. Yet, covering large spatial extents incurs substantial memory and computational overhead. Without a shared global structure, neighboring tiles may drift out of alignment, resulting in inconsistent geometry. Octree hierarchies [12] adaptively allocate capacity to complex regions by subdividing space, but the lack of mechanisms for enforcing consistency across views and scales often causes popping artifacts, missing geometry in under-observed zones, or redundant Gaussians in areas with dense coverage.
In this work, we propose Structure–View Collaborative Gaussian Splatting (SplatCo), a unified framework for high-fidelity reconstruction and rendering of large-scale outdoor scenes. Unlike prior approaches that separately address geometry representation or view consistency, SplatCo integrates both global structure modeling and multi-view coherence into a single, jointly optimized system. We incorporate multi-scale contextual information into a global tri-plane backbone using a three-level compensation mechanism, which aligns coarse scene layout with detailed local geometry and texture. We also introduce a cross-view assisted training strategy that combines the gradient synchronization mechanism proposed in MVGS [13] with a structural-consistency-guided pruning mechanism, which simulatenously governs Gaussian densification and removes overfitted or inconsistent Gaussians based on structural consistency, to ensure consistent updates across overlapping views and stabilize Gaussian positioning and appearance during optimization. By combining these components in a unified training and inference pipeline, SplatCo achieves robust generalization and fine-detail reconstruction across complex, large-scale environments, as shown in Fig. 1.
@@
We validate SplatCo on 13 large-scale outdoor scenes, including four public benchmarks (Mill-19, MatrixCity, Tanks & Temples, WHU) and five custom aerial captures. Compared to state‑of‑the‑art Gaussian‑based methods, SplatCo consistently improves PSNR by 1–2 dB and SSIM by 0.1–0.2, demonstrating its effectiveness in high-fidelity rendering of unbounded outdoor environments.
@@
  • We propose a cross-structure collaboration module that combines global tri‑plane features with local context grids through a hierarchical compensation mechanism, enabling both fine detail preservation and globally consistent geometry.
  • We develop a structural‑consistency‑guided pruning mechanism that removes over‑specialized Gaussians, enhancing geometric integrity and reducing redundancy.
@@
  • We conduct extensive experiments on diverse large‑scale outdoor scenes, showing that SplatCo achieves state-of-the-art performance in detail-preserving novel view synthesis.
II Related Work
II-A Multi-view 3D Reconstruction
@@
3D reconstruction has long been a central problem in computer vision, with many approaches relying on multi-view image inputs to infer scene geometry. Structure-from-Motion (SfM) [14, 15] approaches rely on handcrafted features, such as SIFT [16]), and bundle adjustment [17] to recover sparse scene structures. Multi-View Stereo (MVS) [18, 19] methods extend these techniques by estimating dense depth maps across multiple views, thereby producing detailed point clouds. These raw point clouds can be fed into point orientation methods, such as [20, 21, 22, 23], to compute globally consistent orientations. Meshes can be constructed from the oriented point clouds using surface reconstruction techniques such as Poisson surface reconstruction [24, 25].
@@
Learning-based MVS methods [26, 27, 28, 29, 30, 31] have demonstrated significant improvements in reconstruction accuracy and robustness by adopting a standardized four-stage pipeline: (1) feature extraction, (2) differentiable homography, (3) cost volume regularization, and (4) depth map regression. Variants have improved speed and scalability using recurrent networks [28], coarse-to-fine cascades [29], and Transformer backbones [30]. More compact and flexible depth representations have also been explored [31]. However, cost volumes are memory-intensive and resolution-limited, making these methods less practical for real-time or outdoor-scale reconstruction.
@@
Recently, feature-matching-based approaches, such as DUST3R [32] and its variants [33, 34, 35], further reduce reliance on depth volumes by directly matching features across images. These methods offer a promising alternative for scalable reconstruction.
II-B Novel View Synthesis
@@
Novel view synthesis has become a central task in visual computing, with broad relevance to virtual ad augmented reality, robotics, and and autonomous systems. Early work used explicit representations, such as meshes [36, 37], voxels  [38, 39], and point clouds [40], to model scenes. However, these methods struggled to reproduce fine visual details.
@@
Neural Radiance Fields (NeRF) [1] introduced a view-consistent rendering framework based on implicit volumetric representations learned from 2D images. While NeRF produces realistic results in controlled scenes, its sampling inefficiency and reliance on large MLPs lead to slow training and inference. These issues are especially pronounced in outdoor or large-scale environments. To improve performance, various acceleration methods have been proposed [41, 42], along with anti-aliasing [43, 44], motion deblurring [45, 46], and support for dynamic content [47]. However, the fundamental design of NeRF still poses scalability and efficiency challenges.
@@
3D Gaussian Splatting [7] offers a more efficient rendering pipeline by replacing volume sampling with rasterization and representing the scene as a set of Gaussian primitives. This design enables real-time rendering with competitive visual quality and has become the basis for a growing body of recent work.
II-C 3D Gaussian Splatting for Large-Scale Scenes
@@
3DGS has been widely adopted for modeling objects and scenes due to its speed and rendering quality. Readers may refer to recent surveys for comprehensive overviews of 3DGS and its applications [48, 49, 50]. While early applications focused on compact scenes, such as individual objects [51, 52], humans [53, 54, 55], and indoor environments [56, 57, 58], adapting 3DGS to large-scale outdoor scenes remains a challenge.
@@
To bridge this gap, recent works propose scalable variants of 3DGS. Scaffold-GS [59] introduces a hierarchical, region-aware 3D representation via anchor points. DOGS [60] significantly improves training efficiency through a distributed paradigm that combines spatial partitioning with the ADMM solver. VastGaussian [61] and Hierarchy-GS [62] explore alternative partitioning strategies, each offering distinct architectural advantages. More recent developments include Octree-GS [12], which organizes Gaussian primitives in an octree structure to support adaptive multi-scale rendering, and CityGaussian [63], which incorporates level-of-detail techniques to optimize both training and rendering in urban-scale environments. Inspired by the explicit multi-view training strategy in MVGS [13], we combine it with the multi-level structure–view collaboration mechanism. Our approach requires both structural consistency constraints and cross-view consistency constraints. Computational results demonstrate that this combination is effective for achieving coherent rendering in large-scale unbounded scenes.
@@
Despite these advances, two core limitations remain. First, rigid partitioning of space often breaks continuity of semantic structures, such as roads, facades or vegetation. Second, the lack of structural coordination and view-collaboration mechanisms limits the rendering fidelity required in large-scale scenarios. These challenges often lead to visual artifacts, such as ghosting, texture flickering, and inconsistent shading.
Figure 2: Overview. We propose a Cross-Structure Collaboration Module (CSCM) that hierarchically integrates tri-planes with context grids, effectively preserving fine-grained details while maintaining consistent global geometry. Additionally, we introduce a Cross-View Assisted Training Strategy (CVATS) that seamlessly integrates the Multi-View Augmented Densification (MVAD) and Multi-View Regulated Learning (MVRL) from MVGS [13] with the proposed Cross-View Consistent Loss (CVCL). This synergy enforces across overlapping cameras, effectively mitigating single-view overfitting in large-scale scenes.
III Method
III-A Motivations & Overview
@@
We propose a novel 3D Gaussian Splatting (3DGS) framework that integrates structure-aware and view-aware components to achieve fine-grained scene representation, enabling high-fidelity rendering, especially in large-scale environments. As illustrated in Fig. 2, our framework addresses two key limitations of existing approaches.
@@
First, current 3DGS methods [10, 8, 9] either employ tri-plane structures or use context grid interpolation for spatial modeling. Both strategies have inherent limitations: Tri-plane representations often suffer from detail loss due to fixed-axis-aligned partitioning, while context interpolation lacks global spatial awareness. These issues result in reduced structural expressiveness and visual artifacts.
@@
Second, traditional single-view optimization schemes [7] in 3DGS frameworks exhibit three major drawbacks: 1) susceptibility to overfitting to individual views, 2) discontinuities in regions with low view overlap, and 3) the accumulation of redundant Gaussians near camera paths.
To address the aforementioned limitations, we first introduce a Cross-Structure Collaboration Module (CSCM) that fuses multi-scale tri-plane features with context-aware grid features, enabling the simultaneous preservation of global spatial coherence and local geometric details. We then introduce a unified Cross-View Assisted Training Strategy (CVATS), which includes: 1) multi-view gradient alignment to reduce view-specific overfitting, 2) visibility-aware densification to improve coverage in sparsely observed areas, and 3) adaptive pruning based on geometric consistency to maintain point cloud integrity. Extensive experiments confirm the effectiveness of our proposed method, significantly improving rendering quality across diverse viewing conditions.
III-B Preliminary for 3D Gaussian Splatting

### v4

@@ Abstract
Neural Radiance Fields (NeRFs) have achieved impressive results in novel view synthesis but are less suited for large-scale scene reconstruction due to their reliance on dense per-ray sampling, which limits scalability and efficiency. In contrast, 3D Gaussian Splatting (3DGS) offers a more efficient alternative to computationally intensive volume rendering, enabling faster training and real-time rendering. Although recent efforts have extended 3DGS to large-scale settings, these methods often struggle to balance global structural coherence with local detail fidelity. Crucially, they also suffer from Gaussian redundancy due to a lack of effective geometric constraints, which further leads to rendering artifacts. To address these challenges, we present SplatCo, a structure–view collaborative Gaussian splatting framework for high-fidelity rendering of complex outdoor scenes. SplatCo builds upon three novel components: (1) a Cross-Structure Collaboration Module (CSCM) that combines global tri-plane representations, which capture coarse scene layouts, with local context grid features that represent fine details. This fusion is achieved through the proposed hierarchical compensation mechanism, ensuring both global spatial awareness and local detail preservation; (2) a Cross-View Pruning Mechanism (CVPM) that prunes overfitted or inaccurate Gaussians based on structural consistency, thereby improving storage efficiency while avoiding Gaussian rendering artifacts; (3) a Structure–View Co-learning (SVC) Module that aggregates structural gradients with view gradients, redirecting the Gaussian geometric and appearance attribute optimization more robustly guided by additional structural gradient flow. By combining these key components, SplatCo effectively achieves high-fidelity rendering for large-scale scenes. Comprehensive evaluations on 13 diverse large-scale scenes, including Mill19, MatrixCity, Tanks & Temples, WHU, and custom aerial captures, demonstrate that SplatCo establishes a new benchmark for high-fidelity rendering of large-scale unbounded scenes. Code and project page are available at https://splatco-tech.github.io/.
keywords:
3DGS, real-time rendering, large-scale unbounded scenes, detail preservation, structure-view collaborative learning
1 Introduction
@@
Neural scene representations have significantly advanced novel view synthesis and 3D reconstruction. Among them, Neural Radiance Fields (NeRFs) [relate_nerf] achieve impressive photorealistic rendering of objects and small-scale scenes from multi‑view captures, supporting applications in augmented/virtual reality [intro_vr], autonomous driving [intro_driving], digital twins [intro_twin], telepresence, and cultural heritage preservation [intro_culture].
@@
Applying NeRF to large-scale outdoor and urban environments [exp_meganerf, exp_switchnerf] remains challenging due to its inherent computational and representational limitations. A key difficulty lies in performing differentiable operations over densely sampled 3D points across entire scenes. Even simple operations—such as evaluating a small Multi-Layer Perceptron (MLP) per point—become extremely memory intensive, as all intermediate values must be stored for backpropagation. To overcome these limitations, 3D Gaussian Splatting (3DGS) [relate_3dgs] has emerged as an efficient alternative. It represents a scene as a sparse collection of anisotropic 3D Gaussians whose positions, scales, orientations, and colors are optimized directly from images. These Gaussians are rendered using a tile-based GPU splatting pipeline, enabling high-quality, real‑time rendering.
While 3DGS overcomes the rendering efficiency limitations of NeRFs, its extension to large-scale outdoor scenes  [relate_new_lu2024scaffold, mvgs, liu2024citygaussian] faces three core challenges: global structural coherence, local detail fidelity, and robust cross-view consistency. Current methods address these issues in isolation. Techniques such as Scaffold-GS utilize anchor-based context-grids [method_contextgs, method_hash_grid] to capture fine local details but lack global awareness, while tri-plane representations [method_igs, method_trigs] model the scene’s global features effectively yet lose high-frequency detail information. To balance efficiency and memory, Octree-GS [ren2024octree] adaptively selects appropriate levels from a multi-scale set of Gaussian primitives to ensure consistent rendering performance. However, it fundamentally lacks an effective mechanism to leverage cross-view consistency for accurately identifying and pruning geometrically inaccurate or redundant Gaussians.
Recent promising methods have attempted to address the limitations of conventional 3DGS from multiple perspectives. For example, MVGS [mvgs] introduces a multi-view regulated training strategy to address the limitations of single-view supervision, VastGaussian [lin2024vastgaussian] employs a partition-and-merge strategy to accelerate training convergence, and CityGS-X [citygs-x] proposes a parallelized hybrid hierarchical 3D representation. Nevertheless, these approaches still face the following critical challenges: (1) the absence of a scalable feature encoding scheme for large-scale scenes; (2) the lack of interaction between partitioned blocks leads to inconsistencies at boundaries due to insufficient multi-view constraints; and (3) high memory overhead from maintaining multiple representations, often requiring multiple GPUs and hindering real-time deployment. In practice, however, most users lack access to multi-GPU setups. This raises the question: can we enable large-scale 3DGS training on a single consumer-grade GPU? Therefore, a computationally economical and accessible unified framework that jointly optimizes global structure, local details, and multi-view consistency remains an open problem—and is the primary goal of our work.
In this work, we propose Structure–View Collaborative Gaussian Splatting (SplatCo), for high-fidelity rendering of large-scale outdoor scenes. Unlike prior approaches that primarily focus on supervising view consistency, SplatCo integrates both Structural-View Co-learning (SVC) to achieve robust optimization of Gaussian position and appearance attributes. To robustly extract geometric information from large-scale scenes, we incorporate multi-scale contextual information into a multi-scale tri-plane backbone using a three-level compensation mechanism, achieving a coarse-to-fine cascading optimization. We also introduce a Cross-View Pruning Mechanism (CVPM) that prunes overfitted or inaccurate Gaussians based on structural consistency, thereby improving storage efficiency while avoiding Gaussian rendering artifacts. By combining these components, SplatCo achieves robust fine-detail rendering across complex, large-scale environments, as shown in Fig. 1. SplatCo demonstrates its consistent effectiveness in high-fidelity rendering of unbounded outdoor environments compared to state-of-the-art Gaussian-based methods, and we validate this across 13 large-scale outdoor scenes, including four public benchmarks (Mill-19, MatrixCity, Tanks & Temples, WHU) and five custom aerial captures.
@@
  • We propose a structure–view collaborative Gaussian splatting framework to achieve high-fidelity rendering for large-scale scenes, enabling robust optimization of Gaussian geometry and appearance attributes through aggregating structural gradients with view gradients.
@@
  • We develop a cross-structure collaboration module that combines global tri-plane features with local context grids through the proposed hierarchical compensation mechanism, enabling both fine detail preservation and global spatial awareness.
  • We design a cross-view pruning mechanism that prunes overfitted or inaccurate Gaussians based on structural consistency, thereby improving storage efficiency while avoiding Gaussian rendering artifacts.
@@
  • We conduct extensive experiments on diverse large-scale outdoor scenes with varying structural forms, showing that SplatCo achieves state-of-the-art performance in detail-preserving novel view synthesis.
2 Related Work
2.1 Multi-view 3D Reconstruction
@@
3D reconstruction has long been a central problem in computer vision, with many approaches relying on multi-view image inputs to infer scene geometry. Structure-from-Motion (SfM) [relate_new_cui2017hsfm, relate_new_schonberger2016structure] approaches rely on handcrafted features, such as SIFT [relate_new_zheng2017sift]), and bundle adjustment [relate_new_agarwal2010bundle] to recover sparse scene structures. Multi-View Stereo (MVS) [relate_tra_1, relate_tra_2] methods extend these techniques by estimating dense depth maps across multiple views, thereby producing detailed point clouds. These raw point clouds can be fed into point orientation methods, such as [Hou2022iPSR, Xu2023GCNO, Liu2024DWG, Lin2024WNNC], to compute globally consistent orientations. Meshes can be constructed from the oriented point clouds using surface reconstruction techniques such as Poisson surface reconstruction [Kazhdan2006, Kazhdan2013].
@@
Learning-based MVS methods [relate_mvsnet, relate_mvster, relate_rmvsnet, relate_casmvsnet, relate_transmvsnet, relate_unimvsnet] have demonstrated significant improvements in reconstruction accuracy and robustness by adopting a standardized four-stage pipeline: (1) feature extraction, (2) differentiable homography, (3) cost volume regularization, and (4) depth map regression. Variants have improved speed and scalability using recurrent networks [relate_rmvsnet], coarse-to-fine cascades [relate_casmvsnet], and Transformer backbones [relate_transmvsnet]. More compact and flexible depth representations have also been explored [relate_unimvsnet]. However, cost volumes are memory-intensive and resolution-limited, making these methods less practical for real-time or outdoor-scale reconstruction.
@@
Recently, feature-matching-based approaches, such as DUST3R [relate_dust3r] and its variants [relate_must3r, relate_fast3r, relate_d2ust3r], further reduce reliance on depth volumes by directly matching features across images. These methods offer a promising alternative for scalable reconstruction.
2.2 Novel View Synthesis
@@
Novel view synthesis has become a central task in visual computing, with broad relevance to virtual ad augmented reality, robotics, and and autonomous systems. Early work used explicit representations, such as meshes [relate_new_mesh_liu2019liquid, realet_new_mesh_guo2023vmesh], voxels  [relate_new_voxel_sun2023vgos, relate_new_voxel_schwarz2022voxgraf], and point clouds [relate_new_pointcloud_xiao2023distinguishing], to model scenes. However, these methods struggled to reproduce fine visual details.
@@
NeRFs [relate_nerf] introduce a view-consistent rendering framework based on implicit volumetric representations learned from 2D images. While NeRF methods produce realistic results in controlled scenes, their sampling inefficiency and reliance on large MLPs lead to slow training and inference. These issues are especially pronounced in outdoor or large-scale environments. To improve performance, various acceleration methods have been proposed [relate_new_korhonen2024efficient, relate_new_li2023nerfacc], along with anti-aliasing [relate_mipnerf, relate_new_hu2023multiscale], motion deblurring [relate_deblur_nerf, relate_new_wang2024mp], and support for dynamic content [relate_dnerf]. However, the fundamental design of NeRFs still poses scalability and efficiency challenges.
@@
3D Gaussian Splatting [relate_3dgs] offers a more efficient rendering pipeline by replacing volume sampling with rasterization and representing the scene as a set of Gaussian primitives. This design enables real-time rendering with competitive visual quality and has become the basis for a growing body of recent work.
2.3 3D Gaussian Splatting for Large-Scale Scenes
@@
3DGS has been widely adopted for modeling objects and scenes due to its speed and rendering quality. Readers may refer to recent surveys for comprehensive overviews of 3DGS and its applications [chen2025survey3dgaussiansplatting, fei3dgssurvey, gaosurvey]. While early applications focused on compact scenes, such as individual objects [relate_new_pryadilshchikov2024t, relate_new_sun2024f], humans [relate_new_sakamiya2024avatarperfect, relate_new_chen2025thgs, relate_new_qiu2024anigs], and indoor environments [relate_gsdf, relate_gsrec, relate_gsroom], adapting 3DGS to large-scale outdoor scenes remains a challenge.
@@
To bridge this gap, recent works propose scalable variants of 3DGS. Scaffold-GS [relate_new_lu2024scaffold] introduces a hierarchical, region-aware 3D representation via anchor points. DOGS [chen2024dogs] significantly improves training efficiency through a distributed paradigm that combines spatial partitioning with the ADMM solver. VastGaussian [lin2024vastgaussian] and Hierarchy-GS [kerbl2024hierarchical] explore alternative partitioning strategies, each offering distinct architectural advantages. More recent developments include Octree-GS [ren2024octree], which organizes Gaussian primitives in an octree structure to support adaptive multi-scale rendering, and CityGaussian [liu2024citygaussian], which incorporates level-of-detail techniques to optimize both training and rendering in urban-scale environments. Notably, MVGS [mvgs] introduces a multi-view regulated training strategy to address the limitations of single-view supervision in the original 3DGS, thereby offering partial inspiration for our work.
@@
Despite these advances, two fundamental limitations persist. First, rigid partitioning of space often breaks continuity of semantic structures, such as roads, facades or vegetation. Second, the lack of structural coordination and view-collaboration mechanisms hampers the provision of stable geometric guidance necessary for large-scale scene rendering. These challenges often lead to visual artifacts, such as ghosting, texture flickering, and inconsistent shading.
Figure 2: Overview. we first introduce the CSCM, which integrates multi-scale tri-planes and context-aware grids to generate Hierarchical Dense Embedded features ($f_{h}$), simultaneously ensuring global coherence and local detail preservation. Subsequently, the CVPM maintains integrity by employing a Cross-View Consistency Loss ($L_{CVC}$) and a geometric consistency-based pruning mechanism to reduce redundancy. Finally, the SVC module combines structural and view-dependent gradients for robust joint optimization.
3 Method
3.1 Motivations & Overview
@@
We propose a 3DGS framework tailored for large-scale scenes to achieve high-fidelity scene rendering. As illustrated in Fig. 2, our framework addresses two key limitations of existing approaches:
@@
First, current 3DGS methods [method_contextgs, method_igs, method_trigs] either employ tri-plane structures or use context grid interpolation for spatial modeling. Both technologies have inherent limitations: Tri-plane representations often suffer from detail loss due to fixed-axis-aligned partitioning, while context interpolation lacks global spatial awareness. These issues result in reduced structural expressiveness and visual artifacts.
@@
Second, traditional single-view optimization schemes [relate_3dgs] in 3DGS exhibit two major drawbacks: 1) susceptibility to overfitting to individual views, and 2) the accumulation of redundant Gaussians near camera paths, leading to inefficient storage and training, especially in large-scale scenes.
To address the aforementioned limitations, we first propose a CSCM that fuses multi-scale tri-plane features with context-aware grid features, enabling the simultaneous preservation of global spatial coherence and local geometric details. Second, we design a CVPM, which comprises: 1) a cross-view consistency loss designed to enforce structural coherence across diverse perspectives, and 2) an adaptive pruning strategy governed by geometric consistency to reduce the redundancy of Gaussians. Third, we introduce a SVC module to synergize structural and view gradients, thereby enabling better optimization of Gaussian positions and appearance attributes.
3.2 Preliminary for 3D Gaussian Splatting